six cards are to be drawn or delt, without replacement, from a deck of 52 cards. In how many ways can this be done if
A. all of the cards to be drawn are clubs?
B. All are to be jacks or queens?
C. None are to be aces?
D. All are to be spades or all not spades?

Answer by edjones(7311)

nCr=n!/((n-r)!r!)
.
A) 13C6/52C6=1716 / 20358520=.00008429 (13 clubs)
.
B) 8C6/52C6=28 / 20358520=.000001375 (8 Jacks or Queens)
.
C) 48C6/52C6=12271512 / 20358520=.6028
.
D) I changed this question to what I thought you meant.
(13C6 + 39C6) / 52C6 = (1716 + 3262623)/20358520 =.1603
.
Ed